If one line has a slope of −3 and another has a slope of 2, which line is steeper? Why?
Answer by ewatrrr(24785)


Hi, Line with slope of -3 is the 'steeper line'.
Note:
m = 2/1 is the blue line
m = -3/1 is the green line
m = -5/1 the purple line
In each case, regardless of the 'way' the line slants, the slope tells
us how much y changes as x changes by one unit.
The numerator of the slope reflects 'the change in y'
the larger that numerator is relative to the denominator('the change in x')
The steeper the line slants.